OpenStudy (anonymous):

0.6(y + 3) = 4.8 A. 5 B. 3 C. 13 D. 1.8

OpenStudy (anonymous):

D. 6.0y+1.8=4.8 4.8-1.8=3 3/0.6=1.8

OpenStudy (anonymous):

D